OpenStudy (anonymous):

Find two numbers whose difference is 300 and whose product is a minimum. List them separated with a comma.

OpenStudy (anonymous):

let a = first num let b = sec num let y = min a - b = 300 ab = y a = 300 + b (300+b)b = y b^2 + 300b = y [d/db] [b^2 + 300b] = [d/db][y] dy/db = 2b + 300 dy/db = 0 b = -150 a = 150

OpenStudy (anonymous):

THank you!